# View-Consistent Hierarchical 3D Segmentation Using Ultrametric Feature Fields

### v1

@@ 5.2 View-Consistent Hierarchical Segmentation
| LSeg [26] | 0.218 | 0.406 | 0.670 | 0.432 | 0.698 | 0.536 |
| LSeg + DFF [24] | 0.164 | 0.427 | 0.839 | 0.477 | 0.848 | 0.827 |
| LeRF [19] | 0.266 | 0.413 | 0.857 | 0.512 | 0.598 | 0.641 |
| SAM [22] | 0.443 | 0.652 | 0.787 | 0.627 | 0.808 | 0.462 |
| SAM3D [6] | 0.460 | 0.593 | 0.532 | 0.528 | - | 0.724 |
| Ours, 2D | 0.500 | 0.630 | 0.940 | 0.690 | 1.000 | 0.678 |
| Ours | 0.480 | 0.677 | 0.971 | 0.709 | 1.000 | 0.789 |
@@
| Method | NCfine | NCmid | NCcoar | NCmean |
| LSeg + DFF [24] | 0.224 | 0.485 | 0.706 | 0.4720 |
| SAM3D [6] | 0.510 | 0.608 | 0.668 | 0.596 |
| Ours | 0.499 | 0.633 | 0.828 | 0.653 |
| Ours† | 0.489 | 0.604 | - | - |
| 3D-PIS† [46] | 0.522 | 0.750 | - | - |
@@ 5.4 Implementation Details
We implement our method with help from the publicly available codebase from DFF [24]. We model our feature branch after the RGB branch of Instant-NGP [35] – we employ a multi-resolution grid hash encoder to transform 3D coordinates into features, followed by an MLP. For the grid hash encoder, we configure the number of levels to , features per level to , and the hash map size to . The MLP comprises three hidden layers with dimensions each, producing a final output feature of dimensions.
On Blender-HS and PartNet, we train our model for iterations with a batch size of and use the same optimization parameters as DFF. In contrastive learning (see Eq. 3), we set the temperature to and sample positive and negative pairs from each mask. For depth continuity loss, we sample patches per mask, and begin using the depth continuity loss after iterations. During 3D inference, we extract a point cloud from training-view depth maps, apply voxel downsampling with voxel size , and run outlier removal with distance threshold and number threshold of . We build the graph of points using nearest neighbors, and we transfer point segmentation labels into a novel view using the mode of nearest neighbors. We retain graph components and set the distance threshold to be . Please refer to Sec. 4 for definitions of the above hyperparameters.
@@ 6 Conclusion
For many applications that involve a mobile agent interacting with the real world at scale, such as robotics or augmented reality, a consistent hierarchical segmentation is essential. In this work we have demonstrated significant progress towards that goal, building on state-of-the-art systems that perform multi-granularity object and part detection in images, whose output predictions are neither hierarchical nor consistent across different views. Our ultrametric feature field distills this inconsistent 2D information into a representation that can be queried at will for arbitrary-granularity segmentations that are consistent across views.
7 Acknowledgement
We acknowledge the support of Vannevar Bush Faculty Fellowship and ARL grant W911NF-21-2-0104.
@@ A. Additional Implementation Details
A.1. Hierarchical Sampling
We present the hierarchical sampling algorithm we introduced in Sec. 4.1.3 in Algorithm 1.
A.2. Additional Details on Evaluation Metrics
@@
8.0.1 A.3. Additional Details on Baselines

### v2

@@ 5.2 View-Consistent Hierarchical Segmentation
LSeg [28] | 21.8 | 40.6 | 67.0 | 43.2 | 69.8 | 53.6 |
LSeg + DFF [26] | 16.4 | 42.7 | 83.9 | 47.7 | 84.8 | 82.7 |
LeRF [21] | 26.6 | 41.3 | 85.7 | 51.2 | 59.8 | 64.1 |
SAM [24] | 44.3 | 65.2 | 78.7 | 62.7 | 80.8 | 46.2 |
SAM3D [7] | 46.0 | 59.3 | 53.2 | 52.8 | - | 72.4 |
Ours, 2D | 50.0 | 63.0 | 94.0 | 69.0 | 100.0 | 67.8 |
Ours | 48.0 | 67.7 | 97.1 | 70.9 | 100.0 | 78.9 |
@@ 5.4 Implementation Details
We implement our method with help from the publicly available codebase from DFF [26]. We model our feature branch after the RGB branch of Instant-NGP [37] – we employ a multi-resolution grid hash encoder to transform 3D coordinates into features, followed by an MLP. For the grid hash encoder, we configure the number of levels to , features per level to , and the hash map size to . The MLP comprises three hidden layers with dimensions each, producing a final output feature of dimensions. We provide additional implementation details in the supplementary.
@@
6 Discussion and Limitations
A key limitation of our approach is its dependence on high-quality point clouds produced by the NeRF. While the depth-smoothing loss outlined in Section 5.3 improves point cloud quality, there is much room for further improvement. Regarding our evaluation, we find there is a scarcity of high-quality datasets for hierarchical 3D segmentation that exhibit complex scene structure. Additionally, it is ambiguous which hierarchies are the most meaningful and appropriate in complex environments, without defining end-tasks that rely on these hierarchies. While our Blender-HS Dataset is a first step in providing ground truth on hierarchical 3D segmentation in complex scenes, it is limited to only three scenes. We hope that future efforts can develop better and larger hierarchical 3D datasets, to enable more comprehensive evaluations.
@@ 7 Conclusion
Consistent hierarchical 3D segmentation is essential for many applications involving mobile agents interacting with the real world at scale, such as robotics and augmented reality. In this work we have demonstrated significant progress towards achieving consistent hierarchical segmentation, building on state-of-the-art systems that perform multi-granularity segmentation in images, whose output predictions are neither hierarchical nor consistent across different views. Our ultrametric feature field distills this inconsistent 2D information into a representation that can be queried at will for arbitrary-granularity segmentations that are consistent across views.
Acknowledgements
This work was supported by a Vannevar Bush Faculty Fellowship and ARL grant W911NF-21-2-0104.
@@ A. Additional Implementation Details
8.0.1 A.1. Hyperparameters
On Blender-HS and PartNet, we train our model for iterations with a batch size of and use the same optimization parameters as DFF. In contrastive learning (see Eq. 3), we set the temperature to and sample positive and negative pairs from each mask. We set the loss weight of the Euclidean loss in Eq. 4 to 1. For depth continuity loss, we sample patches per mask, and begin using the depth continuity loss after iterations. During 3D inference, we extract a point cloud from training-view depth maps, apply voxel downsampling with voxel size , and run outlier removal with distance threshold and number threshold of . We build the graph of points using nearest neighbors, and we transfer point segmentation labels into a novel view using the mode of nearest neighbors. We retain graph components and set the distance threshold to be . Please refer to Sec. 4 for definitions of the above hyperparameters.
A.2. Hierarchical Sampling
We first organize the segmentation masks into a hierarchical structure determined by the inclusion ratio between them. One mask is designated as a child of another mask when and . We empirically set and . We present the hierarchical sampling algorithm we introduced in Sec. 4.1.3 in Algorithm 1.
We sample same number of positive pairs and negative pairs for training for training efficiency. Implementing a 4-1 ratio ( more negatives than positives) instead of 1-1, the normalized covering (NC) score increases from 0.709 to 0.720. However, this slow down our training by 43%, primarily due to the time-consuming computation of ultrametric distances and the associated minimum spanning tree.
A.3. Additional Details on Evaluation Metrics
@@
We also evaluate the View Consistency across multiple angles, in Tab. 5. The ranking of the methods is the same.
Method | ||||
LSeg [26] | 0.536 | 0.522 | 0.510 | 0.498 |
LSeg + DFF [24] | 0.827 | 0.813 | 0.810 | 0.808 |
SAM3D [6] | 0.724 | 0.601 | 0.578 | 0.539 |
Ours | 0.789 | 0.763 | 0.742 | 0.712 |
@@
8.0.2 A.4. Additional Details on Baselines
@@
8.0.3 A.5. Training and Inference Time
We train and perform inference on a Titan RTX GPU. Training typically takes 70 minutes, while inference takes 5 seconds per granularity for 10 views. The main expense in inference is the watershed algorithm running on 3D point clouds, which is executed once per granularity and is view-independent.